Math 550 (Section 1)  University of South Carolina
D. Meade  Fall 1997

Homework Problems for Chapter 5 Double and Triple Integrals

Section Page Problems
5.1  300 # 1b^{∗}c, 2b^{∗}c, 4, 6, 8
5.2  313 # 1a, 2c, 3^{∗}, 6^{∗}
5.3  321 # 1ac^{∗}, 2b, 4^{∗}, 7, 8^{∗}, 9, 14
5.4  326 # 1abc^{∗}, 2a^{∗}, 3, 5^{∗}, 9, 12^{∗}, 14
5.6  346 # 1, 4, 5^{∗}, 6, 10, 14, 15^{∗}, 17, 19, 22^{∗}
Review 348 # 1, 2, 3a^{∗}, 6, 8, 9^{∗}, 12, 18^{∗}, 19, 22, 25

\underline{Notes}

• The main purpose of these problems is to provide practice evaluating multiple integrals. For each integral that you evaluate, you should try to decide if the problem could be solved differently, or more easily, by considering a different – but equivalent – formulation of the integral.

• Problems marked with an asterisk (^{∗}) will be collected and graded. In general, problems will be due the Friday following the lecture in which the section was discussed.

• These problems have been selected as the minimum amount of work that you should be doing for this course. Some problems will be straightforward computations; some will require the construction of a logical argument using the terminology and concepts discussed in class; some problems will require you to expand your knowledge (with given information) and apply this to a new setting.

• If you need more practice, there are many other problems in the text that are similar to the ones I have assigned.

• Do not hesitate to ask questions – in class, by e-mail, or during office hours.